Calculate the pH when 200.0 mL of a 1.00 M solution of H2A (Ka1 = 1.0 x 10–6, Ka2 = 1.0 x 10–10) is titrated with the following volumes of 1.00 M NaOH. 250.0 mL of 1.00 M NaOH Group of answer choices
Answer #1


The Balanced Equaton is Inaoht HogA NORA +H2O ginen mole of HgA = MAU=1.0 molx 0.2 L = 0.291 mele of NaoH added a on molx 0.2 and kag - 1.0x1ol, pla q = - lug (kaz) = log (1.0x/0-10) - 10.0 Hendr -San Equation for pH pH = pla Hog CA 2] (HA) = 10 + log